# A Versatile Collage Visualization Technique

### v1

@@ 4 Image-space Collage
In this section, we present our image-space collage methodology, discussing its primary components such as primitive initialization, geometric transformation, and the associated loss functions.
Figure 1 illustrates the framework of our method. Initially, vector primitives, which represent the data, are arranged in a specific initial layout (shall be introduced in Section 4.2). These vector primitives are then rasterized into an image through differential rendering. In this way, we cast the collage optimization problem into the image space, to minimize the rendering loss. Guided by the target image in which the collage is fit, a set of image-based loss functions (discussed in Section 4.3) are computed and back-propagated to a geometric transformation component. Following an update via gradient descent, the parameters within the geometric transformation are optimized to minimize the loss. Subsequently, we apply the updated geometric transformation to the vector primitives. The framework then continues to iterate until a state of minimal loss is attained. The decision to update the geometric transformation, rather than directly modifying the vector primitives, is to maintain the fidelity of data representation in visualization. More details about the geometric transformation process shall be provided in Section 4.1.
@@ 4.1 Geometric Transformation
In collage data visualization, visual primitives encode data via non-positional visual channels such as shape, size, height, or width. It is important to guarantee data encoding fidelity when optimizing the collage visualization. A direct update on the control points of visual primitives in the optimization is unstable, since it is difficult to maintain the original shape or size.
To ensure the positional invariance of visual primitives during the optimization, we introduce geometric transformation into the framework, which includes three operations (i.e., scaling, translation, and rotation) to transform each visual primitive. For each primitive $P_{i}$, its geometric transformation consist of scaling factor $s_{i}$, rotation angle $a_{i}$, and translation offset $(t_{x},t_{y})_{i}$. During the backward propagation process, the control points of primitives remain fixed, and optimization is performed on the parameters related to geometric transformation. With the new geometric transformation, visual primitives are updated for the next iteration.
@@ 4.2 Position Initialization
The initial positioning of visual primitives plays a crucial role in both optimization effectiveness and quality. When primitives are initialized either outside the target shape or at its border, they may be pushed outside the shape due to the primitive overlap loss. Once a primitive completely exits the target shape, it becomes challenging to trace it back solely with loss functions that assess how well the primitives fit within the shape. Escaping primitives can be attracted to moving back into the shape by introducing some distance loss functions, however, this would inevitably increase computational complexity and require more iterations. Similarly, if the initialization positions of primitives are not sufficiently uniformly distributed over the shape, more iterations would be needed, compared to the well-distributed ones.
Proper initial placement of vector primitives facilitates the optimization. In this work, we adopt the Medial Axis Transform (MAT) [15], which can effectively capture the primary contour and shape of objects. By applying the MAT algorithm, we obtain the coordinates of all points on the medial axis and their distances to the nearest boundary point (i.e., the medial width). As Figure 2 shows, with MAT, an area with a larger medial width provides more space to accommodate larger primitives, and vice versa. Given this, we distribute visual primitives along the medial axis, employing a strategy that positions larger primitives at medial points with a greater medial width. By utilizing the coordinates of the medial axis and the medial width for initialization, shape primitives can be distributed more evenly at the skeletal positions inside the shape according to the weights, reducing the number of iterations and improving the filling effect.
Figure 2: MAT-based Position Initialization: with the detected medial axes and their nearest associated widths to the boundary (left), visual primitives are initialized in the way that larger primitives are placed on axes with larger medial widths.
@@ 4.3 Loss Functions
Data Fidelity. In the context of data visualization, it is crucial to maintain data fidelity across non-position visual channels. This means that transformations applied to visual primitives (i.e., scaling, translation, and rotation) should not result in loss of data fidelity. In our framework, there are two strategies to ensure data encoding fidelity. One is to set the scale in geometric transformation as a uniform global value, which will guarantee no distortion in visual encoding. The other strategy is individual scaling for each primitive, instead of applying a global scaling to all primitives. This decision is driven by the observation that global scaling induces changes in every primitive during each iteration, which could lead to instability in the optimization process. The constant shift and fluctuation of primitives under global scaling make it challenging to achieve stable and consistent optimization. To mitigate this, we employ individual scaling that allows us to adjust the size of each primitive independently, yet altering the size of individual primitive. To maintain overall coherence and to ensure that the primitives eventually conform to a unified global scale, we incorporate a Data Fidelity Loss function. Given the original primitive sizes $S_{\text{orig}}=\{s_{\text{orig},i}\}_{i=1}^{n}$ and updated primitive sizes $S_{\text{upd}}=\{s_{\text{upd},i}\}_{i=1}^{n}$, this loss function uses their cosine similarity as a measure:
@@ 7 Summary, limitations and Future Work
In light of the advancements made, there are several promising directions for future research and development. One potential avenue is the exploration of interactive interfaces for target image-space editing in visualization creation. By providing users with intuitive editing and controls in the target image, they can directly manipulate and refine the visual elements in return, allowing for a more interactive and iterative design process. A more concrete avenue can use text-driven editing based on a text-to-image foundation model [9][11].

### v2

@@ 4 Image-space Collage
Our image-space collage methodology consists of three primary components: primitive initialization, geometric transformation, and associated loss functions. Initially, vector primitives representing data are arranged in a specific layout and rasterized into an image through differential rendering. This converts the collage optimization problem into image space to minimize rendering loss. Guided by the target image, image-based loss functions are computed and back-propagated to the geometric transformation component. Using gradient descent, the geometric transformation parameters are optimized to minimize loss, maintaining data representation fidelity. This iterative process continues until minimal loss is achieved.
@@ 4.1 Geometric Transformation
In collage data visualization, ensuring data encoding fidelity is crucial. Directly updating visual primitives is unstable, so we use geometric transformations (scaling, translation, rotation) to maintain shape and size. Each primitive $P_{i}$ has a scaling factor $s_{i}$, rotation angle $a_{i}$, and translation offset $(t_{x},t_{y})_{i}$. During optimization, control points stay fixed while parameters are adjusted. This ensures positional invariance, and the updated geometric transformations are applied iteratively to refine the visualization.
@@ 4.2 Position Initialization
Proper initialization of visual primitives is crucial for effective optimization and quality outcomes in shape filling tasks. Initial placement outside or at the border of a shape risks primitives being pushed out, complicating recovery solely through fitting loss functions. To address this, we employ the Medial Axis Transform (MAT), which identifies primary contours and widths within shapes. Distributing primitives along the MAT facilitates uniform placement, enhancing optimization efficiency and filling quality by leveraging skeletal positions and widths. This approach minimizes iteration needs, improving overall effectiveness in shape filling tasks.
@@ 4.3 Loss Functions
Data Fidelity. Our framework maintains data encoding fidelity through two scaling strategies: global scale for uniformity and individual scaling per primitive to prevent instability. Global scaling affects all primitives uniformly, potentially disrupting optimization. Individual scaling allows independent size adjustments, ensuring stability. To maintain coherence, we employ a Data Fidelity Loss function, measuring cosine similarity between original and updated primitive sizes to guide consistent scaling towards a unified global scale:
@@ 7 Discussion
8 Conclusion and Future Work
In light of the advancements made, there are several promising directions for future research and development. One potential avenue is the exploration of interactive interfaces for target image-space editing in visualization creation. By providing users with intuitive editing and controls in the target image, they can directly manipulate and refine the visual elements in return, allowing for a more interactive and iterative design process. A more concrete avenue can use text-driven editing based on a text-to-image foundation model [9][11]. Also, designing a user interface for collage authoring would be an interesting direction in the future.